OpenStudy (anonymous):

What is the value of (10)C(3)? A. 30 B. 120 C. 720 D. 1000

OpenStudy (anonymous):

ill give medel

OpenStudy (kirbykirby):

\(10 \choose 3\)=\(\frac{10!}{3!\cdot(10-3)!}\) 10! = 10*9*8*7*6*5*4*3*2*1 3!=3*2*1

OpenStudy (anonymous):

help me more plz im confused

OpenStudy (kirbykirby):

\[\frac{10!}{3!\cdot(10-3)!}=\frac{10!}{3!\cdot7!}\] \[=\frac{10\cdot9\cdot8\cdot\underbrace{7\cdot6\cdot5\cdot4\cdot3\cdot2\cdot1}}{(3\cdot2\cdot1)(7\cdot6\cdot5\cdot4\cdot3\cdot2\cdot1)}=\frac{10\cdot9\cdot8\cdot7!}{(3\cdot2\cdot1)7!}=\frac{10\cdot9\cdot8}{3\cdot2\cdot1}=\frac{720}{6}=120\]

OpenStudy (anonymous):

ok thx